10) A lot is in the shape of a right triangle. The shorter leg measures 120 m. The hypotenuse is 40 m longer than the length of the longer leg. How long is the longer leg?
Answer by mananth(16949)

leg 1 = 120 m
longer leg = x
hypotenuse = x+40
(x+40)^2= 120^2+x^2
x^2+80x+1600= 14400+x^2
80x=14400-1600
80x=12800
x= 160 m the longer leg